100 likes | 218 Vues
This report summarizes the muon beam performance for week 26, covering fills 1900 and 1901, and highlights critical statistics such as the number of colliding bunches, integrated luminosity, and average physics rates. Notably, fill 1900 achieved a record in integrated luminosity, while fill 1901 faced issues with online services. Additionally, a technical stop for maintenance and tuning was conducted, focusing on electronic problem studies and noisy thresholds. Observed gaps and system efficiencies are also documented for refining operations moving forward.
E N D
Muon Piquet Report week 26 Oleg Maev, 04.07.2011
Monday, 27.06.2011Fill 1900:Number of bunches per beam = 1237Number of colliding per beam = 1152Integrated Delivered Lumi (1/pb) = 21.20Integrated Recorded Lumi (1/pb) = 19.42Peak Inst Lumi (10^30/cm^2/s) = 379.85Average Inst Lumi (10^30/cm^2/s) = 295.71Average Mu = 1.35Average Pileup = 1.83Average L0 Physics Rate (kHz) = 539.58Average HLT Physics Rate (kHz) = 2.33Efficiency Per Fill (percent) = 90.63Annual Cumulative Efficiency (percent) = 89.44Longest fill and a new record on Integrated LumiRichard played a bit with high Lumi . Random throttling on MUTELL1s has been observed. In General Muon was fine. Only two new tripped Gaps. Oleg Maev, 04.07.2011
Tuesday, 28.06.2011Fill 1901:Number of bunches per beam = 1380Number of colliding per beam = 1296Dumped at ~16:00. == > MD started.Unfortunately, online services (operation plots and Fill summaries) are unavailable still, so I can’t present the numbers.Muon was Ok. Two new tripped Gaps. Oleg Maev, 04.07.2011
MD: Oleg Maev, 04.07.2011
MD: Set of noisy thresholds have been tuned and implemented in the DAQ recipe. Tested without beam. After every physics fill we’ve got more noisy channels. Oleg Maev, 04.07.2011
Technical Stop • OPEN : Monday 4th • UX85B: Survey done by LHCb RP team • expected supervised area: green light to access probably at ~11h00 • END OF DETECTOR WORK: Thur 07 pm • UX85 CLOSURE : From Friday 8 am • Magnet power converter tests at nominal current
Plan and status : 1. RDBs replacement – done 2. MWPC conditioning with negative polarity – prepared 3. Studying of electronic problems – started 4. Global threshold scan - ? TS: Oleg Maev, 04.07.2011
Electronics: Oleg Maev, 04.07.2011
HV status: Oleg Maev, 04.07.2011
Some statistics on new tripped gaps Oleg Maev, 04.07.2011